# Superposition Yields Robust Neural Scaling

### v2

@@ 3 Results
Figure 5: Loss at weak superposition can be well described by the expected number of activated but unlearned features. (a) The actual loss and theory at weak superposition (Equation (4)) agree when weight decay is positive. (b) For those closest to the ideal no superposition case (Appendix D.5), we expect $\alpha_{m}=\alpha-1$, which is close to measured values. Error bars are standard errors.
@@
We begin by analyzing the weak superposition regime. Consider an idealized case where the top $\phi_{1/2}n$ most frequent features are perfectly represented, where $\phi_{1/2}$ is the fraction of represented features. The loss can then be written as:
$L=\sum_{i=1}^{\phi_{1/2}n}\langle(y_{i}-x_{i})^{2}\rangle_{x}+\sum_{i=\phi_{1/2}n+1}^{n}\langle(y_{i}-x_{i})^{2}\rangle_{x}=\sum_{i=\phi_{1/2}n+1}^{n}\langle x_{i}^{2}\rangle_{x}=\langle v^{2}\rangle\sum_{i=\phi_{1/2}n+1}^{n}p_{i}.$ (4)
@@
The first equality follows from the definition of loss. The second holds because the first $\phi_{1/2}n$ features are assumed to be perfectly reconstructed, and the remaining features are ignored (i.e., $y_{i}=0$ for $i>\phi_{1/2}n$). The final expression uses the definition of $x_{i}=u_{i}v_{i}$, where $v\sim U(0,2)$, giving $\langle v^{2}\rangle=4/3$. To approximate the remaining sum, we use the integral $\int_{\phi_{1/2}n}^{n}p_{i}\mathrm{d}i$ and refer to this estimate as the loss by unlearned features. We find that, in the weak superposition regime, the actual losses closely match this prediction (Figure 5a). Slight discrepancies — where actual losses are slightly higher — can be attributed to the fact that models do not exactly learn the first $\phi_{1/2}n$ features. In contrast, in the strong superposition regime, the predicted loss by unlearned features is nearly zero, yet the actual loss remains non-zero, indicating additional sources of error. Focusing on cases closest to the ideal no-superposition scenario — where $\phi_{1/2}n=m$ and the first $m$ features are represented — we observe that such cases occur when $\alpha>1$ and yield a model exponent $\alpha_{m}\approx\alpha-1$ (Figure 5b). This matches the theoretical expectation that $\int_{m}^{n}p_{i}\mathrm{d}i\propto m^{-\alpha+1}$ when $n\gg m$ and $\alpha>1$. Thus, in the weak superposition regime, loss scaling is well described by the contribution of unlearned features, that is, the total frequency of features not represented by the model.
We can now explain when loss follows a power law in the weak superposition regime. When feature frequencies follow a power law, $p_{i}\propto 1/i^{\alpha}$ with $\alpha>1$, the integral $\int_{\phi_{1/2}n}^{n}p_{i}\mathrm{d}i$ becomes a power law in $\phi_{1/2}$. And if the $\phi_{1/2}$ scales approximately as a power of model size $m$ (ideally it is $m/n$). Then the loss also scales as a power law in $m$. Consequently, the deviations from power-law behavior seen in Figure 1b — for non-power-law $p_{i}$ — are consistent with this explanation. Moreover, when $p_{i}\propto 1/i^{\alpha}$ with $\alpha>1$, we expect the model exponent to be $\alpha_{m}\approx\alpha-1$, which matches the relatively small values of $\alpha_{m}$ observed on the right side of Figure 4b. In summary, the weak superposition regime reveals a strong dependence of loss vs. model dimension scaling on the structure of feature frequencies in the data.
Figure 6: Loss scaling at strong superposition is mainly explained via geometry. (a) The more frequent the features are, the more likely their norms are greater than $1$ (purple). (b) The row norm distribution is bimodal around $1$. Related results can be found in Appendix D.6. (c) Fraction of strongly represented features ($\|W_{i}\|_{2}>1$) at stong superposition ($\gamma=-1$) is around $\min\{1,m^{2}/2n\}$ (fast increasing dashed line), which is much larger than $m$ (slowly increasing dashed line). (d) The strongly represented features have decreasing overlaps as expected, where we plotted all the data when $\gamma<0$, and the dashed line is $\kappa^{2}$. (e) At strong superposition, $\alpha_{m}=1$ (horizontal line) if the strongly represented features having small overlaps dominate, which is true for a wide range of $\alpha$. But $\alpha_{m}=2(\alpha-1)$ (fast increasing dashed line) if weakly represented ones with large overlaps dominate, explaining the measured $\alpha_{m}$ when $\alpha$ is large. Error bars are standard errors.
We now turn to the strong superposition regime, where all features are represented. In this regime, the loss arises primarily from the non-zero overlaps between representation vectors. For instance, consider the case where only feature $j$ is activated in the data. Even for $i\neq j$, the output $y_{j}$ receives contributions on the order of $W_{i}\cdot W_{j}$, leading to a loss that scales as $(W_{i}\cdot W_{j})^{2}$ due to the squared error objective. The nonlinearity $\mathrm{ReLU}$ nonlinearity and bias terms do not affect the scaling behavior (Appendix A.1). Consequently, we study the overlaps between representation vectors when the number of vectors exceeds the dimensionality of the space.
We now provide theoretical expectations for vector overlaps. Consider $\nu$ unit vectors $w_{i}\in\mathbb{R}^{m}$ with $\nu\geq m$. It can be shown that [28]
@@
This lower bound, $\kappa$, approaches $\sqrt{1/m}$ when $\nu\gg m$. The bound is met when the vectors form an equal angle tight frame (ETF) [29, 30, 31], which minimizes interference and appears in contexts such as quantum measurements [32, 33, 34, 35] and neural collapse [36, 37]. However, ETFs in real space can only exist if $\nu\leq m(m+1)/2$ [29, 30, 31]. In our toy model, feature frequencies vary, so exact equal-angle configurations are suboptimal — important features may benefit from smaller overlaps. Nonetheless, when feature frequencies are nearly uniform (e.g., small data exponent $\alpha$), $\sqrt{1/m}$ remains a good estimate. Although the rows $W_{i}$ are not strictly unit vectors, they are close enough that the scaling behavior remains valid. In summary, we expect that in the strong superposition regime: (i) Many representation vectors can approximate a mutually orthogonal configuration, with typical overlaps scaling as $\sqrt{1/m}$; (ii) The number of such vectors can be much larger than $m$ but still bounded (e.g., $\sim m^{2}/2$ for ETFs), so when $n$ is too large, some feature vectors cannot be in such configurations, having larger overlaps.
With these theoretical expectations in hand, we evaluate our trained toy models. The first step is to test for the presence of a nearly orthogonal configuration among representation vectors. A preliminary clue comes from examining the norms of the rows $W_{i}$, which exhibit a bimodal distribution: one peak slightly below 1, corresponding to less frequent features, and another above 1, corresponding to the most frequent ones (Figure 6, a and b). This suggests that the most important features are represented in a nearly orthogonal manner, while less important ones exhibit greater overlap. To quantify this, we compute the fraction of rows with norm exceeding 1:
$\phi_{1}=|\{i:\|W_{i}\|_{2}>1\}|/n,$ (6)
which significantly exceeds $m/n$ and is around the ETF expectation $m^{2}/2n$ (Figure 6c). For these rows, the mean squared overlaps match the theoretical bound $\kappa^{2}$ well (Figure 6d). Although these vectors do not form a perfect ETF, their configuration is much closer to an ETF than random initialization (Appendix D.6). In contrast, rows with norms less than $1$ exhibit larger overlaps on average (Appendix D.6). We refer to rows with $\|W_{i}\|_{2}>1$ as strongly represented features and those with smaller norms as weakly represented. These results confirm our expectations: strongly represented features attempt to form an approximately orthogonal configuration with typical overlaps scaling as $\sqrt{1/m}$; their number exceeds $m$ a lot but remains bounded, necessitating the presence of weakly represented features with larger overlaps.
Figure 7: Activation density does not affect scaling exponents in our tests. (a) Loss is roughly proportional to activation density given the degree of superposition ($m=63$, $n=1000$). (b) So, $E$ will only affect the coefficient but not the exponent when considering the power law with model dimension. We plot the evidence $\alpha_{m}\approx 1$ at strong superposition, see more in Appendix D.7.
In the strong superposition regime, the loss is mainly determined by a weighted sum of squared overlaps between representation vectors. For strongly represented features, where overlaps scale as $\sqrt{1/m}$, the loss contribution scales as $1/m$. In contrast, weakly represented features can have large, constant overlaps. In the worst case, their total contribution is proportional to $\sum_{i=\phi_{1}n}^{n}p_{i}\sim m^{-2(\alpha-1)}$, where we approximate $\phi_{1}n\approx m^{2}/2$. (See Appendix A.1 for loss decomposition details) If the loss is dominated by the strongly represented features, we expect the overall loss to scale as $1/m$. In the opposite limit — when weakly represented features dominate — the loss scales as $m^{-2(\alpha-1)}$, meaning the effective model exponent becomes $\alpha_{m}\approx 2(\alpha-1)$. Although we do not solve the toy model analytically, we hypothesize that as $\alpha$ increases (i.e., feature frequencies become more skewed), optimization prioritizes minimizing the loss of the most frequent features. As a result, the remaining loss is mostly due to the weakly represented ones. Empirically, we find a simple and elegant rule (Figure 6e):
$\alpha_{m}\approx\max\{1,2(\alpha-1)\}.$ (7)
In short, the optimization selects whether strongly or weakly represented features dominate the loss, whichever leads to faster decay with model size $m$. We now have a full picture that the origin of $\alpha_{m}=1$ lies in the geometric constraints on vector overlaps, and this scaling is robust as long as the feature frequency distribution is not too skewed.
We also studied the effect of the number of expected activated features or activation density $E$, which was set to $1$. By fixing data exponent $\alpha=1$, which will be shown to be relevant to natural language, we can scan different superposition degrees and activation densities. Since $p_{i}\leq 1$ is required, which is equivalent to $p_{1}\leq 1$, we have $E\leq\sum_{i=1}^{n}1/i^{\alpha}$, setting the upper bound for our scanning. We found that loss is approximately proportional to activation density $E$ (Figure 7a). This fact suggests that the power law exponent should not change, which we confirmed (Figure 7b). Under a controlled superposition degree, activation density linearly increases loss and thus does not affect the scaling exponents in our experiments.
@@
Figure 8: Superposition may explain the neural scaling law observed in actual LLMs. We evaluate four open-sourced model classes, Opt [38], GPT2 [39], Qwen [40], and Pythia [41], which have model sizes from around 100M to 70B (details in Appendix C). (a) We found the mean square overlaps of $W_{i}/\|W_{i}\|_{2}$ roughly follow $1/m$ scaling, where $W$ is the language model head. (b) The model class is reflected by color as panel a, while we use shapes for evaluation datasets [42, 43, 44, 45]. The loss related to model size is fitted as a power law, yielding empirical $\alpha_{m}=0.91\pm 0.04$ close to $1$ (details in Appendix D.8).
@@
Finally, we explore how the phenomena found in our toy model might be relevant to real LLMs [38, 39, 40, 41]. As a simple mapping, we treat tokens as atomic features, with data dimension $n$ equal to the vocabulary size. The model dimension $m$ for LLMs is known. We analyze the language model head, denoted by the weight matrix $W$. Through the norm and interference distributions of the rows of $W$, we claim LLMs are in superposition (Appendix D.8). If we measure token frequency, it follows a power law with exponent $\alpha$ close to $1$ (Appendix D.8). We conclude, based on the knowledge from toy models, that LLMs operate in a superposition regime, and expect loss to be dominated by small overlaps of the strongly represented $\sim\sqrt{1/m}$. We next calculated the mean squared overlaps of normalized rows $W_{i}/\|W_{i}\|_{2}$, and found they roughly obey $1/m$ scaling (Figure 8a). We argue that cross-entropy loss, given that the overlaps are small in absolute value, can be expanded and approximately scales as the mean square overlaps (Appendix A.2). We therefore expect the loss of representation-limited LLMs to have $1/m$ scaling. LLM losses are close to a linear function of $1/m$ (Appendix D.8). Yet when $m\to\infty$, the extrapolation of losses does not hit 00. The non-zero intersection can be due to intrinsic uncertainty in language. Increasing model sizes decreases “wrong" interferences but cannot eliminate uncertainty in the data. So, as in previous papers where loss is decomposed into model size part, dataset size part, and a constant [3], we fit our loss values by the following,
$L=C_{m}/m^{\alpha_{m}}+L_{\backslash m},$ (8)
where the model size part $C_{m}/m^{\alpha_{m}}$ is universal (model size is a function of $m$), and $L_{\backslash m}$ contains loss irrelevant to model size, depending on the evaluation dataset and model class. The fitting yields $\alpha_{m}=0.91\pm 0.04$ (Figure 8b). We inferred from the Chinchilla models [3] that due to model size $N\propto m^{2.52\pm 0.03}$ (Appendix D.8), $\alpha_{m}=(2.52\pm 0.03)\times\alpha_{N}=0.88\pm 0.06$, where $\alpha_{N}=0.35\pm 0.02$ [46] is the power-law exponent of loss with model size. The exponents $\alpha_{m}$ from LLMs are close to 1.
In conclusion, we presented a comprehensive picture through the toy model, at different superposition degrees and data structures, of when the loss scales as a power law with model dimension or size, and what governs the scaling exponent. We confirmed that LLMs operate in the strong superposition regime. Moreover, LLMs [38, 39, 40, 41] quantitatively match toy model predictions: the model exponent $\alpha_{m}$ is close to $1$, and representation overlap scales as $\sqrt{1/m}$. The inferred exponent $\alpha_{m}$ for Chinchilla models [3] is also close to 1. Superposition is therefore an important mechanism underlying the observed scaling laws.
@@ 4 Related works
Neural scaling laws were first characterized empirically [2], demonstrating that for LLMs, the cross-entropy loss improves predictably as a power-law with increased model size (parameters), dataset size, or compute, over multiple orders of magnitude. This finding is built on earlier observations (e.g. [47]) that deep learning performance scales in a smooth power-law fashion with data and model growth. Key works showed the surprisingly universal nature of such scaling behaviors across architectures and tasks [2, 3, 4], directing further development of LLMs.
Several heuristic toy models have been proposed to explain neural scaling laws. One common view is that models aim to fit data manifolds or functions, and the scaling exponents depend strongly on the structure of the data [14, 15]. Another group of models assumes the network learns discrete features or skills [18, 19], whose importance follows a power-law distribution, giving results the same as ours in the weak superposition regime. One toy model predicts that loss scales inversely with model width [24], arguing that parameters independently perform the same task with noise, and the scaling follows from the central limit theorem. However, this model applies mainly in the overparameterized regime.
More formal approaches from statistical learning theory still rely on similar heuristics. The scaling behavior depends on how the limits of dataset size and model size are taken. When the dataset is fixed and model size grows to infinity, the system is variance-limited, and loss scales as $1/m$ by central limit theorem arguments [15]. When the dataset size grows to infinity first, the loss scaling enters the resolution-limited regime. In kernel methods, this leads to $\alpha_{m}=\alpha-1$ [15, 16, 17], consistent with our weak superposition regime. Here, $\alpha$ is the exponent of the power-law decay of kernel eigenvalues, which can be interpreted as abstract feature importance. This regime has also been described as fitting the data manifold [15]. Thus, previous studies mainly address weak superposition or overfitting regimes, while our work reveals new behaviors in the underfitting (where the number of features far exceeds the model dimension) and strong superposition regime, which might be more relevant to LLMs.
Our toy model is Anthropic’s model of superposition [26], with changes in data sampling. The original study explored how data structure influences superposition but did not explicitly control it. Related models have appeared in compressed sensing [48, 49, 50, 51] and neural information processing [52, 53], though key differences arise due to distinct contexts and objectives.

### v3

@@ 3.1 Weak superposition
Figure 4: Loss at weak superposition can be well described by the frequency sum of ignored features. (a) Observation and theory at weak superposition (i.e., Equation (4) as a function of number of represented features, $\phi_{1/2}n$) agree when weight decay $\gamma$ is positive. (b) For those closest to the ideal no superposition case, we expect $\alpha_{m}=\alpha-1$, which is close to measured values. Error bars are standard errors. Details in Appendix D.5.
@@
We seek to understand when loss follows a power law with model dimension, and what determines the exponent when it does in the weak superposition regime. Consider an idealized case where the top $\phi_{1/2}n$ most frequent features are perfectly represented (no overlap), where $\phi_{1/2}$ is the fraction of represented features. The optimal biases are $b_{i}=0$ for $i\leq\phi_{1/2}n$ and $b_{i}=\langle x_{i}\rangle$ for $i>\phi_{1/2}n$ [27]. The loss can then be written as:
$L=\sum_{i>\phi_{1/2}n}\langle(x_{i}-\langle x_{i}\rangle)^{2}\rangle=\sum_{i>\phi_{1/2}n}(\langle v^{2}\rangle p_{i}-\langle v\rangle^{2}p_{i}^{2})\approx\langle v^{2}\rangle\sum_{i>\phi_{1/2}n}p_{i}.$ (4)
@@
The last approximation is right when $p_{i}\ll 1$ for $i>\phi_{1/2}n$ and $p_{i}^{2}$ terms are negligible. We use the definition of $x_{i}=u_{i}v_{i}$, where $v\sim U(0,2)$, giving $\langle v^{2}\rangle=4/3$. We can use the integral $\int_{\phi_{1/2}n}^{n}p_{i}\mathrm{d}i$ to estimate the summation, yielding an expression that depends on the number of represented, $\phi_{1/2}n$, and the data exponent, $\alpha$ (Appendix D.5). We find that, in the weak superposition regime, the actual losses closely match this prediction (Figure 4a). Focusing on cases closest to the ideal no-superposition scenario, where $\phi_{1/2}n=m$ and the first $m$ features are represented, we observe that such cases occur when $\alpha>1$ and yield a model exponent $\alpha_{m}\approx\alpha-1$ (Figure 4b). This matches the theoretical expectation that $\int_{m}^{n}p_{i}\mathrm{d}i\propto m^{-\alpha+1}$ when $n\gg m$ and $\alpha>1$. Thus, in the weak superposition regime, loss scaling is well described by the contribution of unlearned features, that is, the total frequency of features not represented by the model.
We can now answer our Question in the weak superposition regime.
Result 1: “Power law in, power law out" in the weak superposition regime The loss is governed by a sum of frequencies of less frequent and not represented features. Ideally, there are model dimension $m$ most important features being represented. If feature frequencies follow a power law, $p_{i}\propto 1/i^{\alpha}$ with $\alpha>1$, the loss or the summation starting at $m$ will be a power law with $m$ with exponent $\alpha-1$.
This finding of the specific toy model agrees with previous works with very different settings [15, 16, 17, 18, 19, 20], where some power-law skill importance or spectrum is assumed.
3.2 Strong superposition
Figure 5: Loss scaling at strong superposition is explained via geometry. (a) The row norm distribution is bimodal around 11. (b) The more frequent the features are, the more likely their norms are greater than $1$. (c) Variance of squared overlaps for features with $\|W_{i}\|_{2}>1$ is smaller than that of random unit vectors, i.e., $\frac{2(m-1)}{m^{2}(m+2)}$. Overlaps are calculated using directions $W_{i}/\|W_{i}\|_{2}$. We show the measured variances ($\gamma=-0.55$) divided by the above theory value for random vectors. (d) The features with $\|W_{i}\|_{2}>1$ have $1/m$ mean squared overlaps, where we plotted all the data when $\gamma<0$. (e) At strong superposition ($\gamma<0$), $\alpha_{m}=1$ if feature frequencies are flat ($\alpha$ small) due to isotropic vector geometry. But $\alpha_{m}\approx 2(\alpha-1)$ if the feature frequencies are skewed ($\alpha$ large). Error bars are standard errors. More details in Appendix D.6.
We next turn to the strong superposition regime. Consider the case where only feature $j$ is activated. The output $y_{j}$ has activation $\sim W_{i}\cdot W_{j}$, leading to a loss that scales as squared overlaps $(W_{i}\cdot W_{j})^{2}$ due to the definition of loss. The loss arises from the non-zero overlaps between representation vectors. We cannot solve the weight matrix $W$ in this regime. The section goes back and forth between theoretical ansatz and experimental observations to understand the high-level behaviors.
We start by considering relatively even feature frequencies, where trained $W_{i}$ are expected to be isotropic. One simplest theoretical ansatz of isotropic vectors is i.i.d. vectors uniformly on the unit sphere. In $\mathbb{R}^{m}$, the squared overlap of two such random vectors follows $\mathrm{Beta}(\frac{1}{2},\frac{m-1}{2})$ distribution, and therefore has mean $1/m$ and variance $\frac{2(m-1)}{m^{2}(m+2)}\sim 2/m^{2}$. The squared overlaps for isotropic random vectors typically obey $1/m$ scaling.
The actual trained $W_{i}$ have structures whose norms are bimodal near $1$ (Figure 5a), and more important features tend to have larger vector norm (Figure 5b). We want to understand how such a structure will change the scaling of overlaps. It turns out that for better error correction (using bias to cancel interference), the model needs to minimize the maximum overlap rather than the sum of squared overlaps. Consider $\nu$ unit vectors $w_{i}\in\mathbb{R}^{m}$ with $\nu\geq m$. It can be shown that [29]
@@ 3.2 Strong superposition
The lower bound, $\kappa\approx\sqrt{1/m}$ when $\nu\gg m$. The bound is met when the vectors form an equal angle tight frame (ETF) [30, 31, 32], which has no variance in absolute overlaps and appears in contexts such as quantum measurements [33, 34, 35, 36] and neural collapse [37, 38]. ETFs in real spaces can only exist if $\nu\leq\frac{m(m+1)}{2}$ [30, 31, 32]. We find that the $W_{i}$ with $\|W_{i}\|>1$ associated with important features tend to be ETF-like (Figure 5, c and d): the variance of squared overlaps is smaller than that of random vectors and can be near 0 for even feature frequencies (small $\alpha$); the mean of squared overlaps collapse on $1/m\approx\kappa^{2}$. Being ETF or ETF-like can help error correction and reduce loss values, but would not change the typical scaling with $m$ if the number of vectors is much larger than $m$. Similar to ETFs, whose number of vectors is bounded, the number of vectors $W_{i}$ with $\|W_{i}\|>1$ is around $m^{2}/2$ (Appendix D.6), and the less important features tend not to be represented (norm lower than 1). Vectors of these less important features cannot be explained with a simple theoretical ansatz. Yet, combining the lessons from random vectors and ETF, we expect the squared overlaps to scale as $1/m$ robustly for isotropic vectors (confirmed in Appendix D.6). Considering all the overlaps when feature frequencies are even, we then predict the loss to scale as $1/m$, which is true (Figure 5e).
When the feature frequencies are skewed (large $\alpha$), we find that the model exponent $\alpha_{m}$ increases with $\alpha$ and becomes greater than $1$. Vectors being non-isotropic, i.e., important features having much smaller overlaps, may lead to larger $\alpha_{m}$. To illustrate this idea, we conjecture an extreme situation where the $m^{2}/2$ most important features can be ETF-like and contribute negligible loss compared to the less important ones. In the worst case (Appendix A.1), the less important features lead to a loss proportional to $\sum_{i=m^{2}/2}^{n}p_{i}\sim m^{-2(\alpha-1)}$, i.e., $\alpha_{m}=2(\alpha-1)$, which is close to observations (Figure 5e). The real configuration of $W_{i}$ is more complicated than this simple conjecture, requiring advanced future studies on when $\alpha_{m}$ loses robustness and how it depends on feature frequencies sensitively then. To conclude the section, we have
Result 2: Geometric origin of $1/m$ loss scaling ($\alpha_{m}=1$) at strong superposition For even feature frequencies, vectors $W_{i}$ tend to be isotropic in space with squared overlaps scaling like $1/m$ when $n\gg m$, leading to the robust $1/m$ power-law loss. For skewed feature frequencies, representation vectors are heterogeneous in space, making loss sensitive to feature frequencies, where it might need power-law frequencies to have power-law losses.
3.3 LLMs
@@
Figure 6: Superposition may explain the neural scaling law observed in actual LLMs. We evaluate four open-sourced model classes, Opt [39], GPT2 [40], Qwen [41], and Pythia [42], which have model sizes from around 100M to 70B (evaluation details in Appendix C). (a) We found the mean square overlaps of $W_{i}/\|W_{i}\|_{2}$ roughly follow $1/m$ scaling, where $W$ is the language model head. (b) The model class is reflected by color as panel a, while we use shapes for evaluation datasets [43, 44, 45, 46]. The loss related to model size is fitted as a power law, yielding empirical $\alpha_{m}=0.91\pm 0.04$ close to $1$. More analysis in Appendix D.7.
@@
Finally, we explore how our findings might be relevant to real LLMs [39, 40, 41, 42]. As a naive mapping, we treat tokens as atomic features, with data dimension $n$ equal to the vocabulary size. The model dimension $m$ for LLMs is known. We analyze the language model head, denoted by the weight matrix $W$. Through the norm and interference distributions of the rows of $W$, we claim LLMs are in superposition (Appendix D.7). If we measure token frequency, it follows a power law with exponent $\alpha$ close to $1$ (Appendix D.7). We conclude, based on the knowledge from toy models, that LLMs operate in a superposition regime, and expect loss to be related to squared overlaps $\sim 1/m$. We next calculated the mean squared overlaps of normalized rows $W_{i}/\|W_{i}\|_{2}$, and found they roughly obey $1/m$ scaling (Figure 6a). We argue that cross-entropy loss, given that the overlaps are small in absolute value, can be expanded and approximately scales as the mean square overlaps (Appendix A.2). We therefore expect the loss of representation-limited LLMs to have $1/m$ scaling. LLM losses are close to a linear function of $1/m$ (Appendix D.7). Yet when $m\to\infty$, the extrapolation of losses does not hit 0. The non-zero intersection can be due to intrinsic uncertainty in language. Increasing model sizes decreases “wrong" interferences but cannot eliminate uncertainty in the data. So, as in previous papers where loss is decomposed into model size part, dataset size part, and a constant [3], we fit our loss values by the following,
[…]
Result 3: Superposition is an important mechanism behind LLM neural scaling laws LLMs operate in the strong superposition regime. The squared overlaps of token representations scale as $1/m$, token frequencies are flat ($\alpha=1$), and the model size relevant loss scales closely to $1/m$, agreeing with the toy model prediction.
@@ 4 Related works
Neural scaling laws were first characterized empirically [2], demonstrating that for LLMs, the cross-entropy loss improves predictably as a power-law with increased model size (parameters), dataset size, or compute, over multiple orders of magnitude. This finding is built on earlier observations (e.g. [48]) that deep learning performance scales in a smooth power-law fashion with data and model growth. Many works showed the surprisingly universal nature of such scaling behaviors across architectures and tasks [2, 3, 4], directing further development of LLMs.
Several heuristic toy models have been proposed to explain neural scaling laws. One common view is that models aim to fit data manifolds or functions, and the scaling exponents depend strongly on the structure of the data [14, 15]. Another group of models assumes the network learns discrete features or skills [19, 20], whose importance follows a power-law distribution, giving results the same as ours in the weak superposition regime. One toy model predicts that loss scales inversely with model width [25], arguing that parameters independently perform the same task with noise, and the scaling follows from the central limit theorem. However, this model applies in the overparameterized cases and may be less relevant to LLMs.
More formal approaches rely on similar heuristics. The scaling behavior depends on how the dataset size and model size approach infinity. When the dataset is fixed and model size grows to infinity, the system is variance-limited, and loss scales as $1/m$ by central limit theorem arguments [15]. When the dataset size grows to infinity first, the loss scaling enters the resolution-limited regime. In linear models or kernel methods, this leads to $\alpha_{m}=\alpha^{\prime}-1$ [15, 16, 17, 18], seemingly consistent with our weak superposition regime. Here, $\alpha^{\prime}$ is the exponent of the power-law decay of kernel eigenvalues, which can be seen as abstract feature importance. Considering neural tangent kernels, $\alpha^{\prime}$ depends on both data and the model configuration. Our work may be framed as mechanistically showing that $\alpha^{\prime}=\alpha$ ($\alpha$ is the intrinsic data exponent) when models have no superposition, and $\alpha^{\prime}$ is something else when models have strong superposition, which is new. The resolution-limited regime has also been described as fitting the data manifold [15].
Our toy model is based on Anthropic’s model of superposition [27] (an autoencoder), with modifications to the data sampling. The original study explored how data structure influences superposition but did not explicitly control it. Related models have appeared in compressed sensing [49, 50, 51, 52, 53] and neural information processing [54, 55], yet with distinct contexts and objectives. Besides representation, people also studied calculation in superposition [56, 57].